OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 3x – 4y = 8? a) 3/4 b)-4/3 c)-3/4 d)4/3

OpenStudy (kc_kennylau):

Parallel=slope same Perpendicular=slope multiplies to -1

OpenStudy (anonymous):

For two lines to be parallel, they must have the same slope. It will probably help if you first put the original equation into slope intercept form (y = mx + b)

OpenStudy (kc_kennylau):

To put an equation into slope intercept form, solve for y.

OpenStudy (anonymous):

yeah, you're gonna want to rearrange it so that it follows the y=mx+b format, with m being the slope... use what was said above--parallel slope = same slope.. just remember those rules regarding parallel and perpendicular slopes and y=mx+b and you're all good to go :)

OpenStudy (anonymous):

Thank you all so much! :) i wish i could give the both of you medals, because you really helped me out :)